OpenStudy (anonymous):

Fan and Medal!! 273.5 / 20.25 Explain.

OpenStudy (anonymous):

13.50...

OpenStudy (anonymous):

273.5 / 20.25 Both numbers are divisible by 5. So divide them until the lowest fractional form then divide numerator by denominator to get that fractional form into decimal form

OpenStudy (anonymous):

i dont really understand. I was told that I was supposed to make the divisor a whole number. So i have to move 2 spaces to the right... and i get 2025. but when i move 2 spaces to my dividend, i have to add the zero, but i was never told what to do next... @sangya21 plz explain :)

OpenStudy (anonymous):

There are numerous ways to solve this question you can use the method I told above (273.5/5)/(20.25/5) = (54.7/4.05) As neither numerator and denominator have common factor, we will divide 54.7 by 4.05 54.7 / 4.05 = 13.5

OpenStudy (anonymous):

second way is that you make them non decimal number (2735/10)/(2025/100) =2735*100/2025*10 = 27350/2025 (27350/5)/(2025/5) =5470/405 (5470/5)/(405/5) = 1094/81 = 13.5

OpenStudy (anonymous):

oooh. ok! thank you :)

OpenStudy (anonymous):

anytime. :)